ULTRAHIGH-RESOLUTION SPECTROSCOPY OF ANTHRACENE

\underline{MASAAKI BABA}, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan; KAZUTO YOSHIDA, YOSUKE SEMBA, SHUNJI KASAHARA, Molecular Photoscience Research Center, Kobe University, Kobe 657-8501, Japan.

Rotationally resolved fluorescence excitation spectra of the S_{1} ^{1}B_{2u} ← S_{0} transition of anthracene-h_{10} and anthracened_{10} and the change with the magnetic field have been observed in a collimated molecular beam. Recently, the rotational constants in the S_{0} and S_{1} states of anthracene-h_{10} have been reported by Bendkowsky et al. ^{a} We have assigned about 1000 rotational lines and determined more accurate values. We have also performed theoretical calculations and obtained the molecular structure, of which the rotational constants are almost the same as the experimental values. The Zeeman broadening was very small and approximately the same for anthracene-h_{10} and anthracene-d_{10}. It indicates that mixing with the triplet state is very small in the S_{1} ^{1}B_{2u} state. The fluorescence quantum yields of anthracene-h_{10} and anthracene-d_{10} are 0.67 and 0.13, respectively. ^{b} It is concluded that the main nonradiative process in the S_{1} state of the isolated anthracene molecule is not the intesystem crossing to the triplet state, but the internal conversion to the ground state.

aV. Bendkowsky, E. Heinecke and A. Hese, J. Chem. Phys., 127, 224306 (2007)

bM. Sonnenschein, A. Amirav and J. Jortner, J. Phys. Chem., 88, 4214 (1984)